# Complex variational autoencoders admit Kähler structure

### v1

@@ Abstract
It has been discovered that latent-Euclidean variational autoencoders (VAEs) admit, in various capacities, Riemannian structure. We adapt these arguments but for complex VAEs with a complex latent stage. We show that complex VAEs reveal to some level Kähler geometric structure. Our methods will be tailored for decoder geometry. We derive the Fisher information metric in the complex case under a latent complex Gaussian regularization with trivial relation matrix. It is well known from statistical information theory that the Fisher information coincides with the Hessian of the Kullback-Leibler (KL) divergence. Thus, the metric Kähler potential relation is exactly achieved under relative entropy. We propose a Kähler potential derivative of complex Gaussian mixtures that has rough equivalence to the Fisher information metric while still being faithful to the underlying Kähler geometry. Computation of the metric via this potential is efficient, and through our potential, valid as a plurisubharmonic (PSH) function, large scale computational burden of automatic differentiation is displaced to small scale. We show that we can regularize the latent space with decoder geometry, and that we can sample in accordance with a weighted complex volume element. We demonstrate these strategies, at the exchange of sample variation, yield consistently smoother representations and fewer semantic outliers.
@@ 2 Notation and conventions
to denote the complex Hessian with Dolbeault operators, i.e. it is a differential operator on smooth sections. We will use $\Psi$ to denote the (positive) quadratic form in the exponent of a Gaussian density. In general, our notation is mostly standard. For the remainder of our work, we will assume the mixed derivative
@@
are the Wirtinger derivatives and $\mu$ is a mean vector (we are not saying all mixed Hessians vanish). In practice, we will compute Wirtinger derivatives using real-valued surrogates in our VAEs for simplicity reasons. For complex integration, we will use the fact that complex integration has an equivalent form over $\mathbb{R}^{p}$
@@ 5.1 The Fisher information metric in the complex case
$h_{\alpha\overline{\beta}}(z)=4(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)).$ (35)
@@
$\partial_{\alpha}\log p$ $=(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(x-\mu)+(x-\mu)\Sigma^{-1}\partial_{\alpha}\mu-(x-\mu)^{\dagger}\underbrace{\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}}_{=A}(x-\mu)-\underbrace{\text{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma)}_{=t_{\alpha}}$ (41)
$\partial_{\overline{\beta}}\log p$ $=(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}(x-\mu)+(x-\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-(x-\mu)^{\dagger}\underbrace{\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}}_{=B}(x-\mu)-\underbrace{\text{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)}_{=t_{\overline{\beta}}}.$ (42)
@@
$\mathbb{E}[a^{\dagger}Mvv^{\dagger}Nv]=a^{\dagger}M\Sigma Nb.$ (48)
@@
$=\mathbb{E}\Bigg[(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}v(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vv^{\dagger}Bv-(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vt_{\overline{\beta}}$ (50)
$\ \ \ \ \ \ +v\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+v\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-v\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}Bv-v\Sigma^{-1}(\partial_{\alpha}\mu)t_{\overline{\beta}}$ (51)
$\ \ \ \ \ \ -v^{\dagger}Av(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v-v^{\dagger}Avv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+v^{\dagger}Avv^{\dagger}Bv+v^{\dagger}Avt_{\overline{\beta}}$ (52)
$\ \ \ \ \ \ -t_{\alpha}(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v-t_{\alpha}v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+t_{\alpha}v^{\dagger}Bv+t_{\alpha}t_{\overline{\beta}}\Bigg]$ (53)
$=2(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+0+0+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger})+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+0+0+0+0$ (54)
@@
$=2(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger})+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(A\Sigma B\Sigma)$ (56)
$=4(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma\Sigma^{-1}\partial_{\overline{\beta}}\Sigma).$ (57)
This completes the proof. Note that the constants in 1519691 differ.
@@ 5.2 The Fisher information metric and the Hessian of the KL divergence as a Kähler potential
It is well known it statistical theory that the Fisher information is the Hessian of the Kullback-Leibler (KL) divergence. Let the Fisher information with respect to the paramater be defined as
@@
$h_{\alpha\overline{\beta}}=\partial_{\alpha}\partial_{\overline{\beta}}K(z,\overline{z}).$ (61)
@@
$h_{\alpha\overline{\beta}}=\text{Hess}_{z^{\prime},\alpha,\overline{\beta}}\text{KL}(p(x|z)\parallel p(x|z^{\prime})\Bigg|_{z^{\prime}=z}=\partial_{\alpha}\partial_{\overline{\beta}}\text{KL}(p(x|z)\parallel p(x|z^{\prime})\Bigg|_{z^{\prime}=z}=\partial_{\alpha}\partial_{\overline{\beta}}K(z,\overline{z}),$ (62)
@@
subsequently evaluated at $z^{\prime}=z$ once the Hessian is taken.
@@
$\partial\overline{\partial}\ \text{KL}\big(p(x|z)\parallel p(x|z^{\prime})\big)\Bigg|_{z^{\prime}=z}=2(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)),$ (66)
@@
$ds^{2}=d\mu^{\dagger}\Sigma^{-1}d\mu+\text{Tr}(\Sigma^{-1}d\Sigma\Sigma^{-1}d\Sigma).$ (67)
@@
5.3 A log-likelihood sum potential induces a metric asymptotically close to the Fisher information metric
@@
The primary goal of this section is that we will show there exists a Kähler potential that induces a Hermitian metric similarly to the true Fisher information metric. In particular, our method is more justified theoretically rather than simply ”replacing” terms in 35, since if we swap terms for things more computable, it will not necessarily hold that Kähler geometry exists. Thus, this section will attempt to justify we can approximate the Hermitian metric of 35 with something that is valid as a metric that is simultaneously more computable under assumptions.
@@
where we will often take $a_{i}=\frac{1}{N}$. We refer to korotkine2024hessiangaussianmixturelikelihoods for relevant statistical literature on this potential. This Kähler potential is distinct from statistical literature. Taking a logarithm of a mixture sum is not frequent in literature. From a statistical perspective, it poses computational issues rudin2015gmmem. For instance, the sum is in the interior of the logarithm and makes statistical computational difficult, if even relevant. In literature, a similar Kähler potential is studied in xie2023complexrecurrentvariationalautoencoder from a geometry perspective, but without the Gaussian component.
@@
Let us derive this. Define $v_{i}:=x-\mu_{i}$. First, notice, in Einstein notation and since $\mu$ is pluriharmonic,
@@
$\partial_{\alpha}\Psi_{i}=(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\beta}}v_{i}^{\overline{\beta}},\ \ \ \ \ \partial_{\overline{\beta}}\Psi_{i}=(\partial_{\overline{\beta}}v_{i}^{\overline{\delta}})(\Sigma_{i}^{-1})_{\alpha\overline{\delta}}v_{i}^{\alpha}$ (72)
$\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}=(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\overline{\beta}}v_{i}^{\overline{\gamma}})=(\partial_{\alpha}v_{i})^{\dagger}(\Sigma_{i}^{-1})(\partial_{\overline{\beta}}v_{i}).$ (73)
@@
$=\frac{\Omega_{i}}{\rho^{2}}\Big[\frac{1}{\rho^{2}}(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}v_{i}^{\overline{\gamma}}(\partial_{\overline{\beta}}v_{i}^{\overline{\delta}})(\Sigma_{i}^{-1})_{\gamma\overline{\delta}}v_{i}^{\gamma}+(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\overline{\beta}}v_{i}^{\overline{\gamma}})\Big]:=\Xi_{i}.$ (76)
@@
$w_{i}(z)=\frac{a_{i}e^{-\Psi_{i}(z)/\rho^{2}}}{\sum_{j}a_{j}e^{-\Psi_{i}(z)/\rho^{2}}},\ \ \ \ \ \mathbb{E}_{w}[X]=\sum_{i}w_{i}(z)X(z).$ (88)
@@
$h_{\alpha\overline{\beta}}=(1+\frac{1}{\rho^{2}})\mathbb{E}_{w}[\partial_{\alpha}\Psi\partial_{\overline{\beta}}\Psi]-\frac{1}{\rho^{2}}\mathbb{E}_{w}[\partial_{\alpha}\Psi]\mathbb{E}_{w}[\partial_{\overline{\beta}}\Psi]$ (89)
if desired.
@@
We make some other remarks. Suppose we wanted to attempt to approximate the Fisher metric of equation 35 by using discrete substitutions $\mu_{i},\Sigma_{i}^{-1}$. This is nontrivial because attempting to find the closest point among a collection requires a nearest neighbor algorithm. This is especially nontrivial in large-scale settings, which is the focus of this work. This is already baked into the expectation calculation via the softmax function, so our proposed metric of 86, always a valid metric regardless of sampling, has no need for such a nearest-point search algorithm.
@@
$\mathbb{E}_{w}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]=\sum_{i}w_{i}\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}\approx(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu).$ (95)
Proof. The first equality follows from definition of the expectation. For a single Gaussian, notice $w_{i}\approx 1$, $w_{j\neq i}\approx 0$ for $x(z)\approx\mu_{i}$, and by definition
$\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}=(\partial_{\alpha}v_{i})^{\dagger}(\Sigma_{i}^{-1})(\partial_{\overline{\beta}}v_{i}),$ (96)
@@
Figure 2: We show distributional equivalence by plotting the cumulative distribution of the $(1,1)$ real element of the first term Fisher information metric $(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)$ versus its expectation analog $\mathbb{E}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]$ of our proposed metric derivative of the Kähler potential. We sampled 200 points $z$ in the latent space, and we computed the expected value with 300 points. Here, we used the MNIST dataset to train the VAE.
@@
Theorem 2. The potential $K$ we defined in 69 is plurisubharmonic.

### v3

@@ Abstract
It has been discovered that latent-Euclidean variational autoencoders (VAEs) admit, in various capacities, Riemannian structure. We adapt these arguments but for complex VAEs with a complex latent stage. We show that complex VAEs reveal to some level Kähler geometric structure. Our methods will be tailored for decoder geometry. We derive the Fisher information metric in the complex case under a latent complex Gaussian with trivial relation matrix. It is well known from statistical information theory that the Fisher information coincides with the Hessian of the Kullback-Leibler (KL) divergence. Thus, the metric Kähler potential relation is exactly achieved under relative entropy. We propose a Kähler potential derivative of complex Gaussian mixtures that acts as a rough proxy to the Fisher information metric while still being faithful to the underlying Kähler geometry. Computation of the metric via this potential is efficient, and through our potential, valid as a plurisubharmonic (PSH) function, large scale computational burden of automatic differentiation is displaced to small scale. Our methods leverage the law of total covariance to bridge behavior between our potential and the Fisher metric. We show that we can regularize the latent space with decoder geometry, and that we can sample in accordance with a weighted complex volume element. We demonstrate these strategies, at the exchange of sample variation, yield consistently smoother representations and fewer semantic outliers.
@@ 2 Notation and conventions
to denote the complex Hessian with Dolbeault operators, i.e. it is a differential operator on smooth sections. We will denote
$(\partial\overline{\partial}f)_{\alpha\overline{\beta}}=\partial_{\alpha}\partial_{\overline{\beta}}f$ (2)
to be the complex Hessian, since $\partial\overline{\partial}f$ often denotes a $(1,1)$-form. We will use $\Psi$ to denote the (positive) quadratic form in the exponent of a Gaussian density. In general, our notation is mostly standard. For the remainder of our work, we will assume the mixed derivative
@@
are the Wirtinger derivatives and $\mu$ is a mean vector (we are not saying all mixed Hessians vanish). We will show this is empirically supported in Appendix C, Figures 6, 7. In particular, we find our architectures work well with this criterion. We specifically do not restrict our decoder to be holomorphic since it is well known a real-valued function that is holomorphic over a connected and open domain is constant due to the Cauchy-Riemann equations. In practice, we will compute Wirtinger derivatives using real-valued surrogates in our VAEs for simplicity reasons. For complex integration, we will use the fact that complex integration has an equivalent form over $\mathbb{R}^{p}$
@@ 5.1 The Fisher information metric in the complex case
$h_{\alpha\overline{\beta}}(z)=2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)).$ (36)
@@
$\partial_{\alpha}\log p$ $=(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(x-\mu)+(x-\mu)^{\dagger}\Sigma^{-1}\partial_{\alpha}\mu+(x-\mu)^{\dagger}\underbrace{\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}}_{=A}(x-\mu)-\underbrace{\text{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma)}_{=t_{\alpha}}$ (42)
$\partial_{\overline{\beta}}\log p$ $=(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}(x-\mu)+(x-\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+(x-\mu)^{\dagger}\underbrace{\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}}_{=B}(x-\mu)-\underbrace{\text{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)}_{=t_{\overline{\beta}}}.$ (43)
@@
$\mathbb{E}[a^{\dagger}Mvv^{\dagger}Nb]=a^{\dagger}M\Sigma Nb.$ (49)
@@
$=\mathbb{E}\Bigg[(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}v(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vv^{\dagger}Bv-(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}vt_{\overline{\beta}}$ (51)
$\ \ \ \ \ \ +v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+v\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}Bv-v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)t_{\overline{\beta}}$ (52)
$\ \ \ \ \ \ +v^{\dagger}Av(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+v^{\dagger}Avv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+v^{\dagger}Avv^{\dagger}Bv-v^{\dagger}Avt_{\overline{\beta}}$ (53)
$\ \ \ \ \ \ -t_{\alpha}(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v-t_{\alpha}v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-t_{\alpha}v^{\dagger}Bv+t_{\alpha}t_{\overline{\beta}}\Bigg]$ (54)
$=0+(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+0+0+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger})+0+0+0+0+0$ (55)
@@
$=(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger})+\text{Tr}(A\Sigma B\Sigma)$ (57)
$=2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma\Sigma^{-1}\partial_{\overline{\beta}}\Sigma).$ (58)
This completes the proof.
@@ 5.2 The Fisher information metric and the Hessian of the KL divergence as a Kähler potential
It is well known in statistical theory that the Fisher information is the Hessian of the Kullback-Leibler (KL) divergence. Let the Fisher information with respect to the paramater be defined as
@@
$h_{\alpha\overline{\beta}}=\partial_{\alpha}\partial_{\overline{\beta}}K(z,\overline{z}),$ (62)
@@
$h_{\alpha\overline{\beta}}=\text{Hess}_{z^{\prime},\alpha,\overline{\beta}}\text{KL}(p(x|z)\parallel p(x|z^{\prime}))\Bigg|_{z^{\prime}=z}=\partial_{\alpha}\partial_{\overline{\beta}}\text{KL}(p(x|z)\parallel p(x|z^{\prime}))\Bigg|_{z^{\prime}=z}=\partial_{\alpha}\partial_{\overline{\beta}}K(z,\overline{z}),$ (63)
@@
subsequently evaluated at $z^{\prime}=z$ once the (complex) Hessian is taken.
@@
$\partial\overline{\partial}\ \text{KL}\big(p(x|z)\parallel p(x|z^{\prime})\big)\Bigg|_{z^{\prime}=z}=2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)),$ (67)
@@
$ds^{2}=2\text{Re}\ d\mu^{\dagger}\Sigma^{-1}d\mu+\text{Tr}(\Sigma^{-1}d\Sigma\Sigma^{-1}d\Sigma).$ (68)
@@
5.3 A log-likelihood sum potential induces a metric close to the Fisher information metric in the continuum limit
@@
The primary goal of this section is that we will show there exists a Kähler potential that induces a Hermitian metric similarly to the true Fisher information metric. In particular, our method is more justified theoretically rather than simply ”replacing” terms in 36, since if we swap terms for things more computable, it will not necessarily hold that Kähler geometry exists. Thus, this section will attempt to justify we can approximate the Hermitian metric of 36 with something that is valid as a metric that is simultaneously more computable under assumptions.
@@
where we will often take $a_{i}=\frac{1}{N}$. We refer to korotkine2024hessiangaussianmixturelikelihoods for relevant statistical literature on this potential. This Kähler potential is distinct from statistical literature. Taking a logarithm of a mixture sum is not frequent in literature. From a statistical perspective, it poses computational issues rudin2015gmmem. For instance, the sum is in the interior of the logarithm and makes statistical computational difficult, if even relevant. In literature, a similar Kähler potential is studied in xie2023complexrecurrentvariationalautoencoder from a geometry perspective, but without the Gaussian component. We emphasize the positive sign in the above quadratic form, as this matches the theoretical derivation involving $\Psi$ as in Appendix A with
$\Psi\ (\text{here})=\Psi\ (\text{appendix})-\text{other terms}.$ (71)
The derivation with $\Psi$ will not yield the right final signs.
@@
Let us derive this. Define $v_{i}:=x-\mu_{i}$. First, notice, in Einstein notation and since $x$ is pluriharmonic,
@@
$\partial_{\alpha}\Psi_{i}=(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\beta}}v_{i}^{\overline{\beta}}+v_{i}^{\delta}(\Sigma_{i}^{-1})_{\delta\overline{\beta}}(\partial_{\alpha}v_{i}^{\overline{\beta}}),\ \ \ \ \ \partial_{\overline{\beta}}\Psi_{i}=(\partial_{\overline{\beta}}v_{i}^{\overline{\delta}})(\Sigma_{i}^{-1})_{\alpha\overline{\delta}}v_{i}^{\alpha}+v_{i}^{\overline{\delta}}(\Sigma_{i}^{-1})_{\alpha\overline{\delta}}(\partial_{\overline{\beta}}v_{i}^{\alpha})$ (74)
$\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}=(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\overline{\beta}}v_{i}^{\overline{\gamma}})+(\partial_{\overline{\beta}}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\alpha}v_{i}^{\overline{\gamma}}).$ (75)
@@
$=\frac{\Omega_{i}}{\rho^{2}}\Big[\frac{1}{\rho^{2}}[(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}v_{i}^{\overline{\gamma}}+v_{i}^{\delta}(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\alpha}v_{i}^{\overline{\gamma}})][(\partial_{\overline{\beta}}v_{i}^{\overline{\delta}})(\Sigma_{i}^{-1})_{\gamma\overline{\delta}}v_{i}^{\gamma}$ (78)
$\ \ \ \ \ \ \ \ \ +v_{i}^{\overline{\delta}}(\Sigma_{i}^{-1})_{\gamma\overline{\delta}}(\partial_{\overline{\beta}}v_{i}^{\gamma})]+(\partial_{\alpha}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\overline{\beta}}v_{i}^{\overline{\gamma}})+(\partial_{\overline{\beta}}v_{i}^{\delta})(\Sigma_{i}^{-1})_{\delta\overline{\gamma}}(\partial_{\alpha}v_{i}^{\overline{\gamma}})\Big]:=\Xi_{i}.$ (79)
@@
$w_{i}(z)\stackrel{{\scriptstyle!}}{{=}}\frac{a_{i}e^{\Psi_{i}(z)/\rho^{2}}}{\sum_{j}a_{j}e^{\Psi_{j}(z)/\rho^{2}}},\ \ \ \ \ \mathbb{E}_{w}[X]=\sum_{i}w_{i}(z)X(z).$ (91)
@@
$h_{\alpha\overline{\beta}}=(1+\frac{1}{\rho^{2}})\mathbb{E}_{w}[\Phi]-\frac{1}{\rho^{2}}\mathbb{E}_{w}[\partial_{\alpha}\Psi]\mathbb{E}_{w}[\partial_{\overline{\beta}}\Psi]$ (92)
if desired for suitable Hessian surrogate $\Phi$.
It is crucial to understand our derivation for the potential has used $\Psi$. For this reason, we will always work with compact sets around the origin, which is typical in a VAE due to the KL regularization. Note that this is potentially numerically unstable. Thus, an alternative solution is to take
$w_{i}(z)=\frac{a_{i}e^{(\Psi_{i}(z)-M)/\rho^{2}}}{\sum_{j}a_{j}e^{(\Psi_{j}(z)-M)/\rho^{2}}}$ (93)
for suitable $M$, i.e. $M=\max_{i}\Psi_{i}$ for stability reasons.
Note that we reconcile the proof as in here and Appendix A because
$\tilde{h}_{\alpha\overline{\beta}}(z)\propto\mathbb{E}_{w(i)}\left[\mathbb{E}_{x|z,i}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}]+\frac{1}{\rho^{2}}\mathrm{Cov}_{x|z,i}(\partial_{\alpha}\Psi_{i},\partial_{\overline{\beta}}\Psi_{i})\right]+\frac{1}{\rho^{2}}\mathrm{Cov}_{w(i)}\left(\mathbb{E}_{x|z,i}[\partial_{\alpha}\Psi_{i}],\mathbb{E}_{x|z,i}[\partial_{\overline{\beta}}\Psi_{i}]\right),$ (94)
where $\tilde{h}$ is a true target metric, which is an asymptotic identity, thus our methods serves as approximations. We provide empirical evidence of the above through Figure 4. We refer to Appendix B for full details.
@@
We make some other remarks. Suppose we wanted to attempt to approximate the Fisher metric of equation 36 by using discrete substitutions $\mu_{i},\Sigma_{i}^{-1}$. This is nontrivial because attempting to find the closest point among a collection requires a nearest neighbor algorithm. This is especially nontrivial in large-scale settings, which is the focus of this work. This is already baked into the expectation calculation via the softmax function, so our proposed metric of 89, always a valid metric regardless of sampling, has no need for such a nearest-point search algorithm.
@@
$\mathbb{E}_{w}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]=\sum_{i}w_{i}\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}\approx(\partial_{\alpha}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\overline{\beta}}\mu)+(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}\mu)$ (100)
for some $j$.
Proof. The first equality follows from definition of the expectation. For a single Gaussian, notice $w_{j}\approx 1$, $w_{i\neq j}\approx 0$ for $x(z)\approx\mu_{i}$ due to the softmax-type activation, and by definition
$\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}=(\partial_{\alpha}v_{j})^{\dagger}(\Sigma_{j}^{-1})(\partial_{\overline{\beta}}v_{j})+(\partial_{\overline{\beta}}v_{j})^{\dagger}(\Sigma_{j}^{-1})(\partial_{\alpha}v_{j}),$ (101)
@@
Remark. Indeed, we verified empirically the index of the largest $\Psi$ corresponds to the index of the largest weight, which implies $j\neq i$ typically. In particular, the ”winner takes all” in the softmax corresponds to distant $\Psi$.
Figure 2: We show distributional equivalence by plotting the cumulative distribution of the $(1,1)$ real element of the first term Fisher information metric $(\partial_{\alpha}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\overline{\beta}}\mu)+(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}\mu)$ along the maximum index versus computing the expectation of $\mathbb{E}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]$ using the softmax. The purpose of this plot is to convey the asymptotic equivalence of Lemmma 1, i.e. the softmax will be dominated by a singular weight.
@@
Figure 4: We present a cosine similarity between $\operatorname*{arg\,min}_{i}\Psi_{i}$ and $\operatorname*{arg\,max}_{i}\Psi_{i}$ on 50 posterior samples corresponding to $2\times\text{latent dim}$ indices (so one real, one imaginary) of our Kähler metric using pure weights of $w=1$ to replace the softmax function as of 93. The purpose of this plot is to convey the metric elements among these indices are proportional, thus the iterated expectation of 94 holds reasonably well. This figure was created using our MNIST experiment.
@@
Theorem 2. The potential $K$ we defined in 70 is plurisubharmonic.